# Mitigation of multi-path propagation artefacts in acoustic targets with cepstral adaptive filtering

### v1

@@ Abstract
Passive acoustic sensing is a cost-effective solution for monitoring moving targets such as vessels and aircraft, but its performance is hindered by complex propagation effects like multi-path reflections and motion-induced artefacts. Existing filtering techniques do not properly incorporate the characteristics of the environment or account for variability in medium properties, limiting their effectiveness in separating source and reflection components. This paper proposes a method for separating target signals from their reflections in a spectrogram. Temporal filtering is applied to cepstral coefficients using an adaptive band-stop filter, which dynamically adjusts its bandwidth based on the relative intensity of the quefrency components. The method improved the signal-to-noise ratio (SNR), log-spectral distance (LSD), and Itakura–Saito (IS) distance across velocities ranging from 10 to 100 metres per second in aircraft noise with simulated motion. It also enhanced the performance of ship-type classification in underwater tasks by 2.28 and 2.62 Matthews Correlation Coefficient percentage points for the DeepShip and VTUAD v2 datasets, respectively. These results demonstrate the potential of the proposed pipeline to improve acoustic target classification and time-delay estimation in multi-path environments, with future work aimed at amplitude preservation and multi-sensor applications.
@@ I Introduction
The classification and localisation of moving acoustic sources are critical for applications in surveillance, environmental monitoring, ecology, and collision mitigation. Vessels, aircraft, and wildlife animals are common targets of these applications. For those scenarios, passive acoustic sensing offers a cost-effective and non-intrusive solution, especially when using single-sensor setups. However, the sound propagation of moving targets is complex and poses a challenge to such systems. Phenomena such as multi-path reflection, non-linear attenuation, background noise, and frequency shifts are common sources of complexity that obscure the target signal and degrade performance of the solutions [audoly2017measurement, li2024mitigation]. Frequency shifts, often resulting from the Doppler effect characterised by the compression and stretching of the wavefront [rossing2014springer], further contribute to these undesirable distortions.
@@
A central acoustic phenomenon leveraged in these applications is the multi-path reflection caused by interference between direct and surface-reflected sound paths. When a sound-emitting source is in motion, the Lloyd’s Mirror Effect (LME) [jensen2011computational, rodriguez2023fundamentals] is observed through constructive and destructive interference patterns produced by the propagated sound waves. Despite the apparent degradation of the signal, these distinct patterns contain valuable information about the physical characteristics of the target and its motion properties, encoding factors such as velocity, distance, and medium properties. Various applications arise from the study of these patterns, including environmental modelling and acoustic propagation studies [audoly2017measurement], vessel tracking and localisation [ferguson2017convolutional, gaggero2024possibility], marine mammal depth estimation [bouffaut2021source, pereira2020use], and aircraft flight parameter estimation [lo2002aircraft].
@@
Since LME patterns are readily identifiable in spectrograms, the depth of the acoustic source encoded in LME can be estimated when prior knowledge of receiver positioning and expected frequency components are available. In marine ecology, LME patterns are used as references for estimating the vocalisation depth of whales [pereira2020use, bouffaut2021source]. In surveillance applications, the same principles apply to the underwater acoustic localisation of moving targets [zhao2024singlehydrophonebased, meng2016motion] and the estimation of aircraft flight parameters [lo2002aircraft]. However, the presence of this phenomenon can also be undesirable in certain contexts. In the underwater domain, for instance, multi-path reflections are known to contribute to ship collisions with marine mammals, as their perception of sound is disrupted by such interactions [gerstein2005acoustics]. Similarly, these effects increase complexity in classification and recognition tasks within surveillance and defence applications, as the resulting patterns can cause spectral ambiguity, hindering the identification of the true spectral components [domingos2022survey].
@@
Cepstral analysis has proven to be a robust technique for extracting meaningful features from multi-path reflection patterns, as this representation separates the signal and its reflections in an additive manner [oppenheim2004frequency]. The cepstrum is obtained by applying the inverse Fourier transform to the logarithm of the signal’s spectrum. The terminology used to describe cepstral components is adapted from spectral analysis: the counterparts of frequency components are referred to as quefrency, while the counterparts of harmonic components are referred to as rahmonics. Analysis of this representation reveals periodic structures caused by multi-path delays [bogert1963quefrency]. Applications involving cepstral analysis have demonstrated superior performance compared to other techniques, such as correlation analysis, in multi-path environments [gao2008time], which is characteristic of scenarios where LME is present. Additionally, cepstral features are commonly used as input in machine learning applications, enabling accurate range prediction and classification [ferguson2017convolutional, abdul2022mel].
@@
Spectral filtering can be applied to separate multi-path reflection components from acoustic signals [yakubovskiy2015feature, zhao2024singlehydrophonebased]; however, filtering in the cepstral domain presents a strong, and often more robust, alternative. A solution was proposed for time-of-arrival estimation of ship sounds in shallow waters, which involved separating the upper and lower bounds of the cepstrogram, corresponding to low and high quefrencies, from the intermediate quefrency bands. Using the mid-quefrency band, the spectrogram was reconstructed, enabling the effective removal of ship-specific components that were undesirable for the task [ferguson2017convolutional]. A similar strategy was applied for the detection of unmanned aerial vehicles, retaining only the harmonic components associated with target-specific noise [fang2022acoustic]. Both approaches indicate that harmonic components from the target are separable from reflected components through cepstral analysis, although they rely on techniques that remove entire quefrency bands without prior identification of the components.
@@
Improvements to cepstral separation were achieved by averaging the power cepstrum over a sufficiently long time window, providing an estimate of the non-rahmonic components (i.e., those not resulting from multi-path interactions), which could then be subtracted from the original cepstrum. This technique, which was named cepstrum subtraction [ferguson2019improved], has proven effective in enhancing the extraction of rahmonic components for time-delay estimation. Similarly, the removal of cepstral local maxima was applied to reduce multi-path artefacts in hydroacoustic channels [czapiewska2020reduction], improving data transmission quality by up to five times. Both approaches enhance the separation of multi-path components by estimating the behaviour of the rahmonics; however, when applied across the entire quefrency range, they incorporate bands into the filtering process that are not influenced by these artefacts. Moreover, they do not account for the physical constraints of reflection patterns, which could further improve the robustness of such applications.
@@
Despite existing cepstral filtering techniques targeting the separation of reflected components from acoustic data, further improvements can be made by incorporating physical limitations of the environment into the filtering process. Additionally, current methods do not account for variability in medium properties, leaving an opportunity for adaptive filtering of cepstral representations.
@@
In this paper, a temporal filtering approach for spectrograms using cepstral coefficients is proposed, incorporating a band-stop adaptive filter in the cepstral domain. This filter dynamically adjusts its strength based on the relative intensity of each quefrency component, enabling more effective suppression of undesired reflections while preserving target-specific harmonic features. The contributions of this paper are as follows:
@@
    An innovative temporal filtering method for cepstral representations, including the proposal of an adaptive band-stop filter designed to reduce the effects of multi-path propagation in acoustic channels by leveraging the relative intensity of quefrency components.
@@
    A robust evaluation of the proposed filtering method, encompassing both the removal of multi-path reflections in simulated data and its attenuation in a practical classification task.
@@
II Lloyd’s Mirror Effect in Acoustical Recordings
@@
where $A$ is the amplitude of the sound source, $R$ denotes the distance between the source and receiver, and the wave number is given by $k=\frac{2\pi f}{c}$ [rossing2014springer]. In an homogenetic medium bounded with a perfectly reflective surface, the sound waves emitted by the source reflect at the boundaries, and the pressure field measured at the receiver is the combination of the direct and reflected components, as illustrated in Figure 1. This pressure field can be described as
@@
Variations in the distance $r_{r}$ produce a complex interference pattern, known as Lloyd’s Mirror Effect (LME) [jensen2011computational, rodriguez2023fundamentals], resulting from the constructive and destructive interference caused by multi-path reflections of the source wave. The spectrum of a broadband source recorded under these conditions exhibits this complex pattern across all wave numbers, with its characteristics dependent on the properties of the medium. It is important to note that variations in the pattern occur under non-ideal conditions, such as irregular surfaces, multiple boundaries, or stratified media, requiring adaptations to the original formulation in extreme scenarios. In applications where the source information is of interest rather than the medium properties, the LME may represent an undesirable component in the received signal.
@@ III-A Time-filtering of the Cepstrogram
To obtain the cepstral representation, the spectrogram is initially obtained. The Short-Time Fourier Transform (STFT) [rossing2014springer] is commonly used to generate the spectrogram; however, approaches such as Mel frequency spectrograms [stevens1937scale], which produce log-scaled frequencies that mimic human auditory perception, and the Constant Q Transform (CQT) [brown1991calculation], which enhances frequency resolution using geometrically spaced filters, are preferred in domain-specific applications.
Once the time-frequency representation of the signal has been obtained, the cepstral representation needs to be extracted. Following the convention [rossing2014springer], the cepstrum is obtained by performing the inverse Fourier transform of the log-spectrum magnitude, as described by
@@
where $\tau$ denotes the quefrency, $|X(f)|$ is the spectrum magnitude, and $IDFT$ denotes the Inverse Discrete Fourier Transform [rossing2014springer]. Analogously, the cepstrogram is formed by concatenating the cepstra obtained for all spectra in the initial spectrogram.
@@
where $DFT$ denotes the Discrete Fourier Transform [rossing2014springer].
@@ III-B Adaptive Band-Stop Filtering
Despite the simplicity of the frequency filter, real-world recordings often deviate from ideal assumptions, exhibiting variations in target speed, irregularities in the medium boundary, and inhomogeneities within the medium [audoly2017measurement]. In such cases, a fixed corner frequency can be limiting, and adaptive bandwidth selection based on signal features is preferred. As the LME pattern appears in the cepstrogram as quefrency peaks, the presence of reflected components is directly proportional to the cepstral magnitude. Therefore, the bandwidth of the band-stop filter should adapt to amplitude changes, broadening as the amplitude increases and narrowing as it decreases. Thus, when strong multi-path components are present, the filter removes their influence across a wider frequency band.
To implement a filter with adaptive corner frequencies, the variable low-pass filter, initial stage of the motion processing model [brinkworth2009robust], was employed. This filter applies a nonlinearity to the signal and uses the output to select a suitable corner frequency based on a reference range. The process, illustrated in Figure 3, can be divided into four steps.
@@
The first step is the application of a IIR low-pass filter to the signal, following Equation 5, to increase the SNR. By removing high-frequency components, noise is reduced in the representation. Then, to estimate the adaptation state, the result passes through a tone-mapping operation following the Naka-Rushton transform [naka1966spotentials], expressed as
@@ IV-A Acoustic Data with Simulated Movement
In order to evaluate the quality of the filtering pipeline proposed, a set of paired information containing an acoustic signal and its version affected movement artefacts was generated. Spherical spreading behaviour and reflection artefacts were synthetically generated for aircraft machinery sounds, simulating the movement of targets in a bounded media. The machinery sounds were manually selected from the Audio Set dataset [gemmeke2017audioa], a database of manually annotated audio events. The Audio Set dataset contains 10-second sound clips drawn from videos, with a total of 527 target classes. The data chosen for this pipeline was recordings of aircraft in idling state, aiming for sounds produced by targets in stationary state to be used as reference. Given the quality variability and the subjectivity of the labels, the samples were manually selected prioritising sounds without interference from other targets (e.g. human speaking, environmental noise, background music). After the selection, 30 audio segments were selected and resampled to 32000 samples per second to ensure the same frequency range for the spectrograms.
@@
For simulation purposes, the height of the source ($z_{s}$) and the height of the receiver ($z_{r}$) were considered as 100 meters and 0.5 meter above the ground, respectively, with speed of sound in the air being adopted as 343 meters per second [kinsler2000fundamentals]. For each of the aircraft machinery audios present in the dataset, three synthetic versions were produced: one affected by amplitude attenuation only, not considering the time of arrival and, therefore, without delays; one containing only the wave propagated trough the direct path, assuming that there are no reflections; and one with the direct path and reflected components of the audio, simulating the bounded media interaction. For simplicity, the spectrogram of those synthetic versions are denoted in this work as $R(f,t)$, $A(f,t)$ and $B(f,t)$, respectively.
@@
To evaluate the performance of the proposed filtering strategy, $A(f,t)$ and $B(f,t)$ were submitted to the cepstral filtering pipeline. The filtered representations were then compared against $R(f,t)$ using three complementary indicators, aiming to a comprehensive performance evaluation of the filtering approach. The first indicator was the Signal-to-Noise Ratio (SNR), which is used to measure the ratio between the reference spectrogram and noise, considering the latter as the difference between the reference spectrogram and the noisy spectrogram. Thus, the SNR is defined as
@@
and the Itakura-Saito Distance (IS) [itakura1968analysis], defined as
$IS=\sum_{f=1}^{F}{\sum_{t=1}^{T}{\left[\frac{R(f,t)^{2}}{N(f,t)^{2}}-\log\left(\frac{R(f,t)^{2}}{N(f,t)^{2}}\right)-1\right]}},$ (15)
were used as complementary metrics. LSD and IS are metrics designed to measure the spectral distortion of the signal, meaning that lower values are associated with improved quality. By considering the logarithmic compression of the signal, LSD provides a perceptually meaningful evaluation of the distortion between two spectral distributions. It is also less sensitive to amplitude variations and take spectral shape variations in consideration. IS, on the other hand, encompasses both the absolute magnitude and logarithmic ratios, providing a more comprehensive evaluation that resemble the human perception [gray1980distortion]. Both metrics, alongside with the SNR, were used as indicators of the filtering quality in terms of noise attenuation and quality improvement.

### v2

# Mitigation of multi-path propagation artefacts in acoustic targets with adaptive cepstral filtering

@@ Abstract
Passive acoustic sensing is a cost-effective solution for monitoring moving targets such as vessels and aircraft, but its performance is hindered by complex propagation effects like multi-path reflections and motion-induced artefacts. Existing filtering techniques do not properly incorporate the characteristics of the environment or account for variability in medium properties, limiting their effectiveness in separating source and reflection components. This paper proposes a method for separating target signals from their reflections in a spectrogram. Temporal filtering is applied to cepstral coefficients using an adaptive band-stop filter, which dynamically adjusts its bandwidth based on the relative intensity of the quefrency components. The method improved the signal-to-noise ratio (SNR) and log-spectral distance (LSD) across velocities ranging from 10 to 100 metres per second in aircraft noise with simulated motion. It also enhanced the performance of ship-type classification in underwater tasks by 2.28 and 2.62 Matthews Correlation Coefficient percentage points for the DeepShip and VTUAD v2 datasets, respectively. These results demonstrate the potential of the proposed pipeline to improve acoustic target classification and time-delay estimation in multi-path environments, with future work aimed at amplitude preservation and multi-sensor applications.
@@ I Introduction
The classification and localisation of moving acoustic sources are critical for applications in surveillance, environmental monitoring, ecology, and collision mitigation. Vessels, aircraft, and wildlife animals are common targets of these applications [14, 13]. For those scenarios, passive acoustic sensing offers a cost-effective and non-intrusive solution, especially when using single-sensor setups [2]. However, the sound propagation of moving targets is complex and poses a challenge to such systems.
The target signature varies with the movement configuration and surrounding environment, depending on factors such as the distance and grazing angle of approach [38], latitude and longitude effects [26], and the geometry of the environment [41]. These factors cause phenomena that further contribute to undesirable distortions, such as multi-path reflections, non-linear attenuation, and frequency shifts, the latter often resulting from the Doppler effect, characterised by the compression and stretching of the wavefront [37, 3, 27]. When combined with detection or classification tasks, these factors pose a significant challenge to acoustic sensing solutions, highlighting the need for adaptive processing that can adjust its characteristics based on the received stimulus.
@@
A central acoustic phenomenon leveraged in passive acoustic sensing applications is the multi-path reflection caused by interference between direct and surface-reflected sound paths. When a sound-emitting source is in motion, the Lloyd’s Mirror Effect (LME) [23, 36] is observed through constructive and destructive interference patterns produced by the propagated sound waves. Despite the apparent degradation of the signal, these distinct patterns contain valuable information about the physical characteristics of the target and its motion properties, encoding factors such as velocity, distance, and medium properties. Various applications arise from the study of these patterns, including environmental modelling and acoustic propagation studies [3], vessel tracking and localisation [15, 17], marine mammal depth estimation [5, 34], and aircraft flight parameter estimation [29].
@@
Since LME patterns are readily identifiable in spectrograms, the depth of the acoustic source encoded in LME can be estimated when prior knowledge of receiver positioning and expected frequency components are available. In marine ecology, LME patterns are used as references for estimating the vocalisation depth of whales [34, 5]. In surveillance applications, the same principles apply to the underwater acoustic localisation of moving targets [43, 31] and the estimation of aircraft flight parameters [29]. However, the presence of this phenomenon can also be undesirable in certain contexts. In the underwater domain, for instance, multi-path reflections are known to contribute to ship collisions with marine mammals, as their perception of sound is disrupted by such interactions [20]. Similarly, these effects increase complexity in classification and recognition tasks within surveillance and defence applications, as the resulting patterns can cause spectral ambiguity, hindering the identification of the true spectral components [10].
@@
Cepstral analysis has proven to be a robust technique for extracting meaningful features from multi-path reflection patterns, as this representation separates the signal and its reflections in an additive manner [33]. The cepstrum is obtained by applying the inverse Fourier transform to the logarithm of the signal’s spectrum. The terminology used to describe cepstral components is adapted from spectral analysis: the counterparts of frequency components are referred to as quefrency, while the counterparts of harmonic components are referred to as rahmonics. Analysis of this representation reveals periodic structures caused by multi-path delays [4]. Applications involving cepstral analysis have demonstrated superior performance compared to other techniques, such as correlation analysis, in multi-path environments [18], which is characteristic of scenarios where LME is present. Additionally, cepstral features are commonly used as input in machine learning applications, enabling accurate range prediction and classification [15, 1].
@@
Spectral filtering can be applied to separate multi-path reflection components from acoustic signals [42, 43]; however, filtering in the cepstral domain presents a strong, and often more robust, alternative. A solution was proposed for time-of-arrival estimation of ship sounds in shallow waters, which involved separating the upper and lower bounds of the cepstrogram, corresponding to low and high quefrencies, from the intermediate quefrency bands. Using the mid-quefrency band, the spectrogram was reconstructed, enabling the effective removal of ship-specific components that were undesirable for the task [15]. A similar strategy was applied for the detection of unmanned aerial vehicles, retaining only the harmonic components associated with target-specific noise [14]. Both approaches indicate that harmonic components from the target are separable from reflected components through cepstral analysis, although they rely on techniques that remove entire quefrency bands without prior identification of the components.
@@
Improvements to cepstral separation were achieved by averaging the power cepstrum over a sufficiently long time window, providing an estimate of the non-rahmonic components (i.e., those not resulting from multi-path interactions), which could then be subtracted from the original cepstrum. This technique, which was named cepstrum subtraction [16], has proven effective in enhancing the extraction of rahmonic components for time-delay estimation. Similarly, the removal of cepstral local maxima was applied to reduce multi-path artefacts in hydroacoustic channels [9], improving data transmission quality by up to five times. Both approaches enhance the separation of multi-path components by estimating the behaviour of the rahmonics; however, when applied across the entire quefrency range, they incorporate bands into the filtering process that are not influenced by these artefacts. Moreover, they do not account for the physical constraints of reflection patterns, which could further improve the robustness of such applications.
@@
Despite existing cepstral filtering techniques aimed at separating reflected components from acoustic data, further improvements can be achieved by incorporating physical limitations of the environment into the filtering process. Additionally, current methods often do not account for variability in the target signature arising from movement configuration and environment, thereby leaving scope for adaptive filtering of cepstral representations.
@@
In this paper, a temporal filtering approach for spectrograms based on cepstral coefficients is proposed, incorporating an adaptive band-stop filter in the cepstral domain. This filter dynamically adjusts its strength based on the relative intensity of individual quefrency components, enabling more effective suppression of undesired reflections while preserving target-specific harmonic features. The main contributions of this paper are as follows:
@@
    An innovative temporal filtering method for cepstral representations, including an adaptive band-stop filter designed to reduce the effects of multi-path propagation in acoustic channels by leveraging the relative intensity of quefrency components.
@@
    A robust evaluation of the proposed filtering method, encompassing both the removal of multi-path reflections in simulated data and their attenuation in a practical classification task.
@@
II Lloyd’s Mirror Effect in Acoustic Recordings
@@
where $A$ is the amplitude of the sound source, $R$ denotes the distance between the source and receiver, and the wave number is given by $k=\frac{2\pi f}{c}$ [37]. In an homogenetic medium bounded with a perfectly reflective surface, the sound waves emitted by the source reflect at the boundaries, and the pressure field measured at the receiver is the combination of the direct and reflected components, as illustrated in Figure 1. This pressure field can be described as
@@
Variations in the distance $r_{r}$ produce a complex interference pattern, known as Lloyd’s Mirror Effect (LME) [23, 36], resulting from the constructive and destructive interference caused by multi-path reflections of the source wave. The spectrum of a broadband source recorded under these conditions exhibits this complex pattern across all wave numbers, with its characteristics dependent on the properties of the medium. It is important to note that variations in the pattern occur under non-ideal conditions, such as irregular surfaces, multiple boundaries, or stratified media, requiring adaptations to the original formulation in extreme scenarios. In applications where the source information is of interest rather than the medium properties, the LME may represent an undesirable component in the received signal.
@@ III-A Time-filtering of the Cepstrogram
To obtain the cepstral representation, the spectrogram is initially obtained. The Short-Time Fourier Transform (STFT) [37] is commonly used to generate the spectrogram; however, approaches such as Mel frequency spectrograms [39], which produce log-scaled frequencies that mimic human auditory perception, and the Constant Q Transform (CQT) [7], which enhances frequency resolution using geometrically spaced filters, are preferred in domain-specific applications.
Once the time-frequency representation of the signal has been obtained, the cepstral representation needs to be extracted. Following the convention [37], the cepstrum is obtained by performing the inverse Fourier transform of the log-spectrum magnitude, as described by
@@
where $\tau$ denotes the quefrency, $|X(f)|$ is the spectrum magnitude, and $IDFT$ denotes the Inverse Discrete Fourier Transform [37]. Analogously, the cepstrogram is formed by concatenating the cepstra obtained for all spectra in the initial spectrogram. It is worth noting that this formulation corresponds to the real cepstrum, which is derived from the log-magnitude spectrum and therefore does not retain phase information. While phase information can encode fine temporal structure, magnitude-based cepstral representations are known to effectively capture spectral periodicities associated with harmonics and reflections [35].
@@
where $DFT$ denotes the Discrete Fourier Transform [37].
@@ III-B Adaptive Band-Stop Filtering
Despite the simplicity of the frequency filter, real-world recordings often deviate from ideal assumptions, exhibiting variations in target speed, irregularities in the medium boundary, and inhomogeneities within the medium [3]. In such cases, a fixed corner frequency can be limiting, and adaptive bandwidth selection based on signal features is preferred. As the LME pattern appears in the cepstrogram as quefrency peaks, the presence of reflected components is directly proportional to the cepstral magnitude. Therefore, the bandwidth of the band-stop filter should adapt to amplitude changes, broadening as the amplitude increases and narrowing as it decreases. Thus, when strong multi-path components are present, the filter removes their influence across a wider frequency band.
To implement a filter with adaptive corner frequencies, the variable low-pass filter, initial stage of the motion processing model [6], was employed. This filter applies a nonlinearity to the signal and uses the output to select a suitable corner frequency based on a reference range. The process, illustrated in Figure 3, can be divided into four steps.
@@
The first step is the application of an IIR low-pass filter to the signal, following Equation 5, to increase the SNR. By removing high-frequency components, noise is reduced in the representation. Then, to estimate the adaptation state, the result passes through a tone-mapping operation following the Naka-Rushton transform [32], expressed as
@@ IV-A Acoustic Data with Simulated Movement
In order to evaluate the quality of the filtering pipeline proposed, a set of paired information containing an acoustic signal and its version affected movement artefacts was generated. Spherical spreading behaviour and reflection artefacts were synthetically generated for aircraft machinery sounds, simulating the movement of targets in a bounded media. The machinery sounds were manually selected from the Audio Set dataset [19], a database of manually annotated audio events. The Audio Set dataset contains 10-second sound clips drawn from videos, with a total of 527 target classes. The data chosen for this pipeline was recordings of aircraft in idling state, aiming for sounds produced by targets in stationary state to be used as reference. Given the quality variability and the subjectivity of the labels, the samples were manually selected prioritising sounds without interference from other targets (e.g. human speaking, environmental noise, background music). After the selection, 30 audio segments were selected and resampled to 32000 samples per second to ensure the same frequency range for the spectrograms.
@@
For simulation purposes, the height of the source ($z_{s}$) and the height of the receiver ($z_{r}$) were considered as 100 meters and 0.5 meter above the ground, respectively, with speed of sound in the air being adopted as 343 meters per second [25]. For each of the aircraft machinery audios present in the dataset, three synthetic versions were produced: one affected by amplitude attenuation only, not considering the time of arrival and, therefore, without delays; one containing only the wave propagated trough the direct path, assuming that there are no reflections; and one with the direct path and reflected components of the audio, simulating the bounded media interaction. For simplicity, the spectrogram of those synthetic versions are denoted in this work as $R(f,t)$, $A(f,t)$ and $B(f,t)$, respectively.
@@
To evaluate the performance of the proposed filtering strategy, $A(f,t)$ and $B(f,t)$ were submitted to the cepstral filtering pipeline. The filtered representations were then compared against $R(f,t)$ using two complementary indicators, aiming to a comprehensive performance evaluation of the filtering approach. The first indicator was the Signal-to-Noise Ratio (SNR), which is used to measure the ratio between the reference spectrogram and noise, considering the latter as the difference between the reference spectrogram and the noisy spectrogram. Thus, the SNR is defined as
@@
was used as complementary metric. LSD is designed to measure the spectral distortion of the signal, meaning that lower values are associated with improved quality. By considering the logarithmic compression of the signal, LSD provides a perceptually meaningful evaluation of the distortion between two spectral distributions. It is also less sensitive to amplitude variations and take spectral shape variations in consideration [21]. Both LSD and SNR were used as indicators of the filtering quality in terms of noise attenuation and quality improvement.